# Xiangqi Wang1, Xiuying Chen2, Mohamed Elhoseiny3, Xiangliang Zhang1,†

### v1

@@ 3.3 Image Generation By Self-Validation
In the self-validation process $\mathcal{F}$, for each image $\mathcal{I}_{ij}^{d}$, based on its image description, $\mathcal{M}_{v}$ is prompted to generate a set of simple questions $\Phi_{ij}^{d}=\{\phi_{ij1}^{d},\phi_{ij2}^{d},\dots,\phi_{ijp}^{d}\}$ (e.g.,“Is there a wooden chair in the image?”), where $p$ denotes the question number to evaluate the alignment. The function $\mathcal{F}$ takes the image $\mathcal{I}_{ij}^{d}$, its description $\mathcal{T}_{ij}^{d}$, and the set of questions $\Phi_{ij}^{d}$ as inputs and outputs an alignment score $S_{ij}^{d}$, which is calculated as the ratio of correctly answered questions to the total number of questions:
$S_{ij}^{d}=\mathcal{F}(\mathcal{I}_{ij}^{d},\mathcal{T}_{ij}^{d},\Phi_{ij}^{d})$
We set a threshold $\zeta$, where: (i) If $S_{ij}^{d}<\zeta$, the image $\mathcal{I}_{ij}$ will be reworked in line with the description until it meets the required standard; (ii) If $\zeta\leq S_{ij}^{d}<1$, the image meets the basic criteria but contains an error $\mathcal{E}_{ij}^{d}$, which will be documented; and (iii) If $S_{ij}^{d}=1$, the image is considered to fully align with the description and is deemed acceptable.
3.4 Test Case Generation & Evaluation
Q&A generation with error control. To enhance the accuracy of question generation, particularly when addressing potential flaws in images, we propose error control. Despite thorough self-validation, it’s not guaranteed that every image will be flawless. Furthermore, when generating problems, we aim to avoid introducing biases stemming from the visual capabilities of examiner LVLM (Zhang et al., 2024b). Therefore, when generating questions, we will only include the image description $\mathcal{T}_{ij}^{d}$ and any identified defects $\mathcal{E}_{ij}^{d}$ in the input to the examiner $\mathcal{M}_{v}$. The function $\mathcal{M}_{v}$ generates the question $Q_{ij}^{d}$ based on the image description and errors:
$Q_{ij}^{d}=\mathcal{M}_{v}(\mathcal{T}_{ij}^{d},\mathcal{E}_{ij}^{d}).$
This will enable the creation of a diverse set of questions, along with reference answers, that specifically target the defective elements. For each image, we will provide a related question $Q_{ij}^{d}$ (e.g., multiple-choice or true/false). These questions, along with the accompanying images, will be presented to the LVLMs under evaluation for their response.
Evaluation. The response $\mathcal{P}_{ij}^{d}$ from the tested LVLMs was compared to the reference answer $A_{ij}^{d}$ to determine accuracy. If $\mathcal{P}_{ij}^{d}$ matched $A_{ij}^{d}$, it was marked correct ($\text{Acc}_{ij}^{d}=1$); otherwise, it was marked incorrect ($\text{Acc}_{ij}^{d}=0$). The overall accuracy $\text{Acc}^{d}_{\text{total}}$ was calculated as the average accuracy over all $N$ questions.
4 Experiment
In this section, we evaluate seven of the latest models using Autobench-V and perform human evaluations to validate our experimental findings. First, we demonstrate how Autobench-V significantly reduces potential answer leakage and self-enhancement bias, as evidenced by the experimental results in Figure 5. Next, based on Figure 6, Table 1, Table 2, and LABEL:tab:average_scores_transposed, we analyze the impact of various evaluation factors (e.g., user input) and question difficulty on model performance, which reveals several insightful findings. We then present the model rankings across five user input categories with varying difficulty levels in Figure 8, followed by a discussion on human evaluations regarding alignment during Autobench-V’s generation process. Lastly, we investigate position bias in the evaluation process, as illustrated in Figure 9.
@@ 4 Experiment
4.1 Experimental Setup
@@
Selected models. In evaluating LVLMs, we selected seven representative models: GPT-4o, GPT-4o mini (Achiam et al., 2023), Claude-3.5-Sonnet, Claude-3-Haiku (Anthropic, 2024), Gemini-1.5-Flash (DeepMind, 2024), GLM-4v (GLM et al., 2024), and the open-source Qwen2-VL (Bai et al., 2023a), detailed in Table 7. These advanced models exhibit exceptional image understanding. Some well-known open-source models, such as Llava-1.6 (Liu et al., 2024a) and MiniGPT-4 (Zhu et al., 2023), were tested and found to perform poorly. Additionally, their capabilities differ significantly from other models, so they are not discussed in our evaluation. We chose GPT-4o as the examiner model for generating image descriptions, questions, and answers due to its strong overall performance. The descriptions were then passed to Flux-pro (blackforestlabs, 2024), a text-to-image model known for outstanding image generation. We also experimented with other text-to-image models (Rombach et al., 2022; Podell et al., 2023; Betker et al., 2023). However, their performance was suboptimal.This combination enables effective automated generation of image-based questions, crucial for the evaluation process (Ying et al., 2024; Fu et al., 2024; Xu et al., 2023).
@@
Experimental setting. We set $n=4$ for the number of general aspects and $m=6$ for the number of fine-grained aspects, as this configuration yields the highest diversity in the generated aspects as illustrated in Table 1, allowing for a broader range of scenes and elements. We set $\omega$ = 10, namely 10 pictures for each fine-grained aspect. Therefore, we evaluate 720 images for each user input (with each user input having three difficulty levels). For easy difficulty, we set the self-validation threshold $\zeta_{e}=1$ since the scenes are simpler and contain fewer elements, which justifies a higher threshold. For the medium and hard difficulty levels, the images contain more elements, so we lower the thresholds to $\zeta_{m}=\zeta_{h}=0.8$ to avoid compromising efficiency. The error control mechanism that follows ensures the appropriateness of this threshold.
@@
Table 2: Average performance (Accuracy) of all models at different difficulty levels.
@@
GPT-4o 90.43% 79.81% 75.02%
GPT-4o mini 88.01% 76.98% 70.70%
Gemini-1.5-Flash 88.07% 74.64% 70.85%
Claude-3.5-Sonnet 89.28% 75.49% 63.67%
Claude-3-Haiku 86.82% 73.40% 67.42%
GLM-4v 90.43% 77.29% 64.93%
Qwen2-VL 89.57% 79.03% 71.89%
Figure 5: Performance of various LVLMs across three difficulty levels when provided only with image descriptions, without the corresponding images.
Figure 6: Overall performance of different models at easy, medium, and hard levels.
4.2 Examiner Priority
To mitigate the potential for answer leakage associated with self-enhancement bias (e.g., when the model being evaluated is also utilized for generating the evaluation cases) in the examiner LLM, we enhance the fairness of the assessment by having AutoBench-v generate questions from image descriptions rather than directly from images. This approach separates the visual information from the generation process, reduceing the risk of self-enhancement bias (Ye et al., 2024) that could occur if questions were derived from the examiner model’s (GPT-4o) visual capabilities, which might cause unfair comparison. By employing only textual descriptions for generation, we eliminate the influence of GPT-4o’s specific visual processing abilities, thereby ensuring a more equitable evaluation.
To validate the fairness of this method, we conducted an experiment in which models were presented with image descriptions alongside corresponding questions while avoiding direct access to the images. As demonstrated in Figure 5, the results revealed consistent performance across all models, with minimal variance (0.4% for easy questions and 2.4% for hard questions). This consistency suggests that, in the absence of visual input, models’ textual understanding ability is almost equal, which means the benchmark effectively assesses visual comprehension and does not show obvious bias towards the examiner LVLM (i.e., GPT-4o).
Figure 7: Score variation of models from easy to hard across different user inputs. As task difficulty increases, the performance disparity between models becomes more pronounced.
4.3 Main Results
As shown in Figure 6, through the evaluation of various models on AutoBench-V, we can observe several findings that can bring insights for future work. More detailed results are in Table 6.
Table 3: Average accuracy for various user inputs at different difficulty levels.
User Input Easy Medium Hard
Basic. 90.59% 75.90% 63.33%
Spatial. 82.46% 69.14% 63.00%
Seman. 91.28% 79.84% 74.52%
Reason. 86.50% 74.67% 68.97%
Atmos. 94.76% 83.33% 75.66%
Model performance decreases as task difficulty increases, with GPT-4o showing the strongest average performance across tasks. This trend is consistent across all models, with scores steadily declining as the difficulty increases from easy to hard, as shown in Figure 13. For example, GPT-4o’s average score drops from 90.43% at the easy level to 75.02% at the hard level. Despite the overall decline, GPT-4o maintains its leading position across all difficulty levels. Additionally, the result highlights that the most notable shift occurs between easy and medium. Although a few samples show improved scores with increased difficulty mildly, the majority trend still experiences a decline, reinforcing the validity of our difficulty grading mechanism.
As task difficulty increases, the performance disparity between models becomes more pronounced. As illustrated in Figure 7, the performance decline across models varies with increasing task difficulty. Models like Claude-3.5-Sonnet and GLM-4v experience more pronounced drops. In difficult semantic tasks, GPT-4o maintains a strong score of 79.36%, compared to Claude-3.5-Sonnet and GLM-4v, which achieve 66.97% and 66.82%, respectively, highlighting GPT-4o’s superior ability to handle complex abstractions. The standard deviations across models at the three difficulty levels are 1.26%, 2.14%, and 3.74%, indicating increased disparity as task difficulty rises. Notably, models like GPT-4o and GLM-4v show more consistent performance with smaller score variations, suggesting stability across difficulty levels. In contrast, models like Claude-3.5 and Qwen2-VL exhibit greater score fluctuations, indicating higher sensitivity to difficulty changes.
Models demonstrate superior performance in semantic and atmospheric understanding while lagging in spatial and reasoning tasks. As illustrated in LABEL:tab:average_scores_transposed, our results reveal a consistent pattern across difficulty levels. Models excel in semantic and atmospheric understanding, maintaining high accuracy even at the hard level of 74.52% and 75.66% respectively. In contrast, spatial and reasoning prove more challenging, with accuracy dropping to 63.00% at the hard level. These findings indicate that while LVLMs have developed strong capabilities in comprehending semantic content and scene atmosphere, they still struggle with tasks involving spatial relationships and complex visual reasoning. To address these limitations, we suggest future research could explore training strategies that emphasize enhancing spatial reasoning and complex visual problem-solving capabilities in LVLMs.
4.4 Model Rank Overview
@@
Figure 8 reveals distinct performance patterns among different models under various difficulty. Notably, models like GPT-4o, while not exhibiting a significant advantage in simpler tasks, demonstrate outstanding performance in more challenging scenarios (e.g.,hard questions). Conversely, models such as GLM-4v perform well on easier tasks but show diminished capabilities as task difficulty increases. This indicates an imbalance in the model’s capabilities across different difficulty levels, highlighting the importance of cross-difficulty evaluation. It also demonstrates that AutoBench-V is effective in revealing such imbalances.

### v3

@@ 3.3 Image Generation By Self-Validation
In the self-validation process $\mathcal{F}$, for each image $\mathcal{I}_{ij}^{d}$, based on its image description, $\mathcal{M}_{v}$ is asked to generate a set of simple questions $\Phi_{ij}^{d}=\{\phi_{ij1}^{d},\phi_{ij2}^{d},\dots,\phi_{ijp}^{d}\}$ (e.g.,“Is there a wooden chair in the image?”), where $p$ denotes the question number to evaluate alignment. The function $\mathcal{F}$ takes the image $\mathcal{I}_{ij}^{d}$, its description $\mathcal{T}_{ij}^{d}$, and the set of questions $\Phi_{ij}^{d}$ as inputs and outputs an alignment score $S_{ij}^{d}$, which is calculated as the ratio of correctly answered questions to the total number of questions, denoted as $S_{ij}^{d}=\mathcal{F}(\mathcal{I}_{ij}^{d},\mathcal{T}_{ij}^{d},\Phi_{ij}^{d})$. We set a threshold $\zeta$, where: (i) If $S_{ij}^{d}<\zeta$, the image $\mathcal{I}_{ij}$ will be reworked in line with the description until it meets the required standard; (ii) If $\zeta\leq S_{ij}^{d}<1$, the image meets the basic criteria but contains an error $\mathcal{E}_{ij}^{d}$, which will be documented for further processing (described in subsection 3.4); and (iii) If $S_{ij}^{d}=1$, the image is considered to fully align with the description and is deemed acceptable.
@@
Figure 4: Models’ accuracy (here we take “basic understanding” as an example) before and after the option adjustment, without image input (only guess the answer by textual information).
3.4 Test Case Generation & Evaluation
Q&A generation with error control. To enhance question generation accuracy, especially for addressing potential image flaws, we propose an error control mechanism. While self-validation helps, not all images can be guaranteed flawless. To mitigate biases from relying on a single examiner LVLM (Zhang et al., 2024b), we incorporate a diverse set of examiners, inspired by prior work Bai et al. (2024). This approach reduces self-enhancement bias Ye et al. (2024) and promotes greater diversity in evaluation.
When generating questions, we will only include the image description $\mathcal{T}_{ij}^{d}$ and any identified defects $\mathcal{E}_{ij}^{d}$ in the input to the examiners $\mathcal{M}_{v}^{k}$, where $k$ represents the index of the selected examiner from the set of available examiners. Each time a question is generated, an examiner $\mathcal{M}_{v}^{k}$ will be randomly selected. The function $\mathcal{M}_{v}^{k}$ generates the question $Q_{ij}^{d}$ based on the image description and errors, denoted as $Q_{ij}^{d}=\mathcal{M}_{v}^{k}(\mathcal{T}_{ij}^{d},\mathcal{E}_{ij}^{d})$.
Moreover, the options for each question will also be generated by a randomly selected examiner from the set. For each image, we will provide a related question $Q_{ij}^{d}$ (e.g., multiple-choice or true/false). These questions, along with the accompanying images, will be presented to the LVLMs under evaluation for their response.
Error-driven option adjustment. A key challenge in evaluating LVLMs is the generation of distractor options that are both plausible and challenging. Distractors generated by LVLMs are often trivial, leading to inflated performance, even under high-difficulty settings or purely text-based evaluations, as models can infer correct answers without relying on visual inputs (Chen et al., 2024b). Attempts to directly create harder options often result in low-quality, overly obscure distractors.
To address this, we propose error-driven option adjustment, a mechanism to improve distractor difficulty while preserving quality. For a given question $Q_{ij}^{d}$, with correct option $O_{ij}^{c}$ and distractors $O_{ij}^{f}=\{O_{ij,1}^{f},\dots,O_{ij,k}^{f}\}$. First, the randomly selected LVLM examiner is given $Q_{ij}^{d}$ and options $\{O_{ij}^{c}\}\cup O_{ij}^{f}$, but the correct option $O_{ij}^{c}$ is intentionally mislabeled as incorrect. The model generates a plausible “correct answer” $O_{ij}^{f*}$, given by $O_{ij}^{f*}=\mathcal{M}_{v}^{k}(Q_{ij}^{d},O_{ij}^{c})$. This generated $O_{ij}^{f*}$ replaces one or more original distractors in $O_{ij}^{f}$, resulting in an updated set $O_{ij}^{f^{\prime}}$. This adjustment increases the difficulty of distinguishing the correct answer from distractors, as $O_{ij}^{f^{\prime}}$ now contains plausible alternatives generated by the model itself under adversarial conditions. Evaluation. The response $\mathcal{P}_{ij}^{d}$ from the tested LVLMs was compared to the reference answer $A_{ij}^{d}$ to determine accuracy. The overall accuracy was calculated as the percentage of correctly answered questions. By given the reference answer, we use LLM-as-a-Judge Zheng et al. (2023) to label the responses.
4 Experiments
4.1 Experiment Setup
Table 2: Performance (accuracy) of all models on five user inputs and three difficulty levels (red and blue represent the highest and lowest scores, respectively, in each row). Claude-3.5 is Claude-3.5-Sonnet and Llama-3.2-90B is Llama-3.2-90B-Vision.
User Input Difficulty Model
GPT-4o GPT-4o mini Gemini-1.5-Flash Gemini-1.5-Pro Claude-3.5 Claude-3-Haiku GLM-4v-Plus Qwen2-VL Llama-3.2-90B
Basic. Easy 77.03% 72.73% 78.47% 78.47% 84.69% 70.81% 74.64% 76.56% 50.72%
Medium 56.65% 46.35% 44.21% 51.93% 57.51% 44.64% 47.64% 53.88% 32.19%
Hard 41.15% 30.53% 34.07% 39.82% 45.58% 31.42% 34.07% 45.58% 29.78%
Spatial. Easy 59.22% 44.69% 60.00% 60.56% 65.00% 42.78% 57.22% 62.78% 36.31%
Medium 34.50% 22.71% 29.69% 41.48% 38.86% 25.33% 31.44% 36.68% 24.45%
Hard 21.46% 15.14% 23.32% 27.35% 26.01% 15.84% 17.49% 25.91% 14.55%
Seman. Easy 73.10% 64.62% 68.53% 70.56% 72.59% 55.84% 62.94% 66.33% 42.64%
Medium 61.90% 46.75% 54.11% 58.44% 61.90% 46.75% 51.95% 61.04% 39.57%
Hard 43.64% 33.79% 44.55% 47.27% 45.45% 26.36% 36.36% 44.55% 30.14%
Reason. Easy 57.06% 46.20% 56.40% 57.56% 57.00% 36.63% 46.51% 53.49% 36.26%
Medium 47.16% 35.22% 40.43% 47.39% 46.09% 25.22% 36.96% 48.03% 26.64%
Hard 38.21% 30.66% 21.43% 35.71% 36.59% 14.29% 24.53% 39.15% 25.00%
Atmos. Easy 60.20% 55.94% 58.91% 66.83% 60.40% 48.02% 52.48% 56.72% 36.63%
Medium 39.73% 30.67% 39.21% 39.65% 44.93% 27.75% 33.48% 39.21% 28.63%
Hard 33.79% 18.26% 28.64% 30.45% 34.55% 15.00% 23.18% 31.05% 22.73%
Average 49.65% 39.63% 45.46% 50.23% 51.81% 35.11% 42.06% 49.40% 31.75%
@@ 4.1 Experiment Setup
Models. In evaluating LVLMs, we selected nine representative models: GPT-4o, GPT-4o mini (Achiam et al., 2023), Claude-3.5-Sonnet, Claude-3-Haiku (Anthropic, 2024), Gemini-1.5-Flash (DeepMind, 2024a), Gemini-1.5-Pro (DeepMind, 2024b), GLM-4v-Plus (GLM et al., 2024), and the open-source models: Qwen2-VL (Bai et al., 2023), Llama-3.2-90B-Vision (Meta, 2024), detailed in Table 5. These advanced models exhibit exceptional image understanding. Some well-known open-source models, such as Llava-1.6 (Liu et al., 2024) and MiniGPT-4 (Zhu et al., 2023a), were tested and found to perform poorly and exclude finally from the model list. We include GPT-4o, Gemini-1.5-Pro, and Claude-3.5-Sonnet in the examiner pool as examiner models for generation due to their strong overall performance. The descriptions were then passed to Flux-1.1-Pro (Labs, 2023), a text-to-image model known for outstanding image generation. We also experimented with other text-to-image models (Rombach et al., 2022; Podell et al., 2023; Betker et al., 2023) but did not include them because of their suboptimal performance.
@@
Hyperparameter Setting. We set $n=4$ for the number of general aspects and $m=6$ for the number of fine-grained aspects, as this configuration yields the highest diversity in the generated aspects as illustrated in Table 1, allowing for a broader range of scenes and elements. We set $\omega$ = 10, namely 10 pictures for each fine-grained aspect. Therefore, we evaluate 720 images for each user input (with each user input having three difficulty levels). For easy difficulty, we set the self-validation threshold $\zeta_{e}=1$ since the scenes are simpler and contain fewer elements, which justifies a higher threshold. For the medium and hard difficulty levels, the images contain more elements, so we lower the thresholds to $\zeta_{m}=\zeta_{h}=0.8$ to avoid compromising efficiency. It balances minimizing errors and avoiding excessive image rejection; a smaller threshold risks inaccuracies, while a larger one may discard too many images. The error control mechanism ensures high-quality question generation.
Table 3: Average accuracy for various user inputs at different difficulty levels (red and blue represent the highest and lowest scores, respectively, in each column).
User Input Easy Medium Hard Average
Basic. 73.79% 48.33% 36.88% 53.00%
Spatial. 54.28% 31.68% 20.78% 35.58%
Seman. 64.13% 53.60% 39.12% 52.28%
Reason. 49.67% 39.24% 29.51% 39.47%
Atmos. 55.12% 35.92% 26.41% 39.15%
4.2 Main Results
Performance drops significantly with increasing difficulty. As shown in LABEL:tab:combined_scores and LABEL:tab:average_scores_transposed, across all models and user inputs, scores decrease as the difficulty progresses from easy to medium to hard. In the basic understanding, the average score across all models drops from approximately 73.79% at easy to 36.88% at hard. This trend is consistent across the LABEL:tab:combined_scores, where the hard is noticeably shorter than the easy and medium.
Basic understanding is the simplest task, while spatial understanding represents the most challenging one: Most models perform better in basic understanding compared to spatial understanding. For example, the average score for basic understanding at easy difficulty is 73.79%, while spatial understanding achieves lower averages of 54.28%. This indicates the imbalance of current LVLMs’ capabilities and emphasizes the importance of holistic improvement.
Some models achieve scores below 25% on tasks with medium and hard difficulty, which suggests that the model performs worse than random guessing. As shown in LABEL:tab:combined_scores, GPT-4o Mini, Gemini-1.5-Flash, Claude-3.5-Haiku, GLM-4v-Plus, and Llama-3.2-90B-Vision achieved scores below 25% on hard tasks involving spatial understanding and reasoning capacity. This indicates that AutoBench-V effectively challenges the evaluated models, mitigating the answer leakage issue where models rely solely on textual information to answer correctly (we will discuss this in detail in subsection 4.3).
@@ 4.2 Main Results
Table 4: Average performance (Accuracy) of all models at different difficulty levels (red and blue represent the highest and lowest scores, respectively, in each column).
@@
GPT-4o 65.32% 47.99% 35.65%
GPT-4o mini 56.83% 36.34% 25.68%
Gemini-1.5-Flash 64.46% 41.53% 30.40%
Gemini-1.5-Pro 66.79% 47.78% 36.12%
Claude-3.5-Sonnet 67.93% 49.86% 37.63%
Claude-3-Haiku 50.82% 33.94% 20.58%
GLM-4v-Plus 58.76% 40.29% 27.13%
Qwen2-VL 63.17% 47.77% 37.25%
Llama-3.2-90B-Vision 40.51% 30.30% 24.44%
4.3 Exploratory Experiments
We aim to develop a deeper understanding of the following questions at this section:
  • Q1: What impact does error-driven option adjustment have on visual question answering?
  • Q2: What are the advantages of multi-evaluator assessment compared to single-evaluator evaluation?
  • Q3: Does the position of correct options affect model performance?
  • Q4: Is the ranking of the same model consistent across different user inputs and different levels of difficulty?
Figure 5: Comparison of answer distribution under position bias conditions. Correct answers at A v.s. correct answers are evenly distributed at A,B,C,D.
Unnecessary visual information was significantly reduced after the option adjustment. As illustrated in Figure 4, in the absence of image input, the scores of various models before option adjustment were concentrated within the range of 40 to 60. Interestingly, without option adjustment, the “guess accuracy” significantly increases with the difficulty level. For instance, GPT-4o can guess less than 40% answers correctly on the easy level while achieving a guessing accuracy of 60% on the hard level. After option adjustment, the performance of all models declined significantly, demonstrating that error-driven option adjustment effectively mitigates answer leakage in the questions.
Figure 6: The score variance of GPT-4o, Gemini-1.5-Pro, and Claude-3.5-Sonnet when using different models as examiners.
Under the setting of multiple examiners, the performance differences among models become more pronounced. As shown in Figure 6, when multiple evaluators are employed, the variance in the performance of the three models is significantly larger. This suggests that using a single model for evaluation may reduce the difference gap between models, leading to an “evaluation bottleneck” while the multi-examiner can mitigate this to achieve a more diverse assessment and lead to a larger performance gap across models.
@@ 4.3 Exploratory Experiments
The position of the correct options can influence model performance. To avoid potential position bias in the evaluation process, we ensured an even distribution of correct options. To further explore the possible effects of position bias on model performance, we conducted a case study where all correct answers were intentionally placed in option A and compared it to the evenly distributed case (25% per option) to assess the necessity of maintaining even distribution, as shown in Figure 5. The deviation rate was calculated using the formula $R=\frac{S_{A}-S_{U}}{S_{U}}$, where $S_{A}$ is the score with correct answers concentrated in option A, and $S_{U}$ is the score under evenly distributed conditions. As illustrated in Figure 5, most models exhibit decreased performance after the answer distribution is balanced, indicating a strong preference for selecting option A.
Different models have distinct performance patterns under various difficulties. As shown in Figure 7, Claude-3.5-Sonnet achieves the best performance, consistently ranking highest across most tasks, while Llama-3.2-90B-Vision performs the weakest, frequently ranking last. Interestingly, models like Qwen2-VL rank higher on hard tasks than on easy ones, whereas models like Gemini-1.5-Pro excel on easy tasks but decline slightly on hard ones. These findings highlight the importance of difficulty-based evaluation and demonstrate AutoBench-V’s effectiveness in revealing such performance imbalances.